OpenStudy (anonymous):

If a:b = 3:5 and a:c = 5:7, then what is the value of b:c? Please Help And Please Explain!

OpenStudy (gw2011):

I think b:c=7:9 because a in the a:b is 3 and then a in the a:c increases by 2 to get 5. So, if you look at a:b and a:c the difference is 2.

OpenStudy (anonymous):

Ah Thank you!:) i really appreciate it:)

OpenStudy (gw2011):

You're welcome

OpenStudy (anonymous):

its 25:21. a/b=3/5 so b/a=5/3 and a/c=5/7 so multiplying (b/a)(a/c) gives you b/c which is (5x5)/(7x3)=25/21

OpenStudy (anonymous):

That makes More sense Thank You:) I Really Appreciate it:)